OpenStudy (mathhater):

Need Help!! Will FAN and MEDAL!!! 1)Little Timmy wants to measure how tall his house is. He doesn’t have a tape measure but does have a stopwatch. He recruits Big Bob to climb on his roof and drop a rock. He times the rocks fall and finds it to be 1.36 s. How tall is his house?

OpenStudy (osprey):

this is the KINEMATIC EQUATIONS again. s = ut + 1/2 a t squared. u is zero "dropped". s= 1/2 x10 x 1.36^2=about 1/2 x 10 x 2 = about 10metres ? If I'm right, that is. http://perendis.webs.com